1.A gas occupies a volume of 400cm³ at 500k and atmosphere pressure.what will be the temperature of the gas when the volume and pressure of the gas is 100cm³ and 0.5atomsphere respectively
2.Gas A takes 10sec to diffuse through a distance of one fifth of a meter.Another gas B takes the same time to diffuse through a distance of 10cm.If the rfm of gas A is 60, calculate the molecular mass of gas B
1
Expert's answer
2020-04-28T12:26:13-0400

1. pV=nRT

n=pV/RT= 1 x 0.4 /0.082 x 500 = 0.01

T=pV/nR = 0.5 x 0.1 / 0.01 x 0.082 = 62.5 K.

2.